# Reasoning or Simply Next Token Prediction? A Benchmark for Stress-Testing Large Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive quantitative performance on a wide range of benchmarks, natural language processing [20, 17], general knowledge question-answering[10, 5], and coding [3, 13]. Additionally, by integrating with some advanced prompting techniques, such as Chain-of-Thought (CoT) [18] and its variants [19, 16, 21], LLMs seem to exhibit a certain level of reasoning abilities in solving question-answer problems. However, some studies [12] have raised concerns about data leakage (i.e., training models on the test sets), potentially rendering these results unreliable. These seemingly contradictory findings prompt the question of whether LLMs are genuinely performing reasoning tasks or merely predicting the next token. If LLMs are truly capable of reasoning, they should remain unaffected by the replacement of key symbols within the test set.
A hallmark of human intelligence is the ability to handle abstract concepts and to associate them with arbitrary terms [14]. With a few exceptions such as onomatopoeia, the connection between particular words and particular meanings is arbitrary, and identical concepts are invoked by different words in different human languages (e.g. dog vs chien). Similarly, human reasoners are capable of analogizing structural relationships from one domain to another, meaning that conceptual equivalence can be retained even when details change [9]. It follows that true human-like comprehension should be unimpaired when terms are substituted for synonymous terms, as long as the substitution is comprehensibly defined.
We wondered whether LLM peformance reflects true human-like comprehension in this sense, or whether it relies heavily on the specific terms used on training corpora. To assess this, we propose MMLU-SR, a new benchmark dataset that uses symbol replacement to remove some important terms from the questions and answers as shown in Fig. 1. Instead of relying on memorized terms, this approach tests whether LLMs can reason using the definitions and concepts of those terms, ensuring a more robust evaluation of their understanding.
Our evaluations on gpt-3.5-turbo, llama3-8b, and gemini-1.0-pro showed significantly lower performance on MMLU-SR compared to the original MMLU, demonstrating the effectiveness of our approach in preventing models from exploiting memorized data. MMLU-SR thus provides a more challenging and revealing test of LLMs’ true reasoning abilities and understanding.
@@
2 Related Work
MMLU Variants Benchmarks. MMLU Variants such as CMMLU [11] and TMMLU+ [15] are adaptations of the MMLU benchmark for non-English languages; they translate the original MMLU questions and answers into other languages, providing a way to evaluate language models’ performance in non-English contexts. These benchmarks are crucial for assessing the generalizability and robustness of models across different languages and cultural settings. They preserve the original structure and intent of MMLU while enabling a broader examination of multilingual capabilities.
Reasoning Capabilities Benchmarks. Several advanced reasoning benchmarks have been developed to evaluate the reasoning capabilities of language models. AGIEval [22] includes standardized tests and civil service exams to assess reasoning and problem-solving skills in academic and professional scenarios. BoolQ [4] comprises over 15,000 real yes/no questions paired with Wikipedia passages to test the ability of models to infer answers from contextual information. GSM8K [6] features 8.5K grade-school math problems requiring multistep operations, targeting the evaluation of basic to intermediate mathematical problem-solving abilities. DROP [7], an adversarially-created reading comprehension benchmark, challenges models to navigate references and perform discrete operations such as addition and sorting, thus evaluating their capacity to understand complex texts and execute logical reasoning tasks.
@@ 3.1 Dataset Construction
We have developed the MMLU-SR benchmark to rigorously evaluate the reasoning and understanding capabilities of LLMs. Inspired by ObjectNet [2], our benchmark contains three subsets: “Question Only”, “Answer Only”, and “Question and Answer”, each offering a unique perspective on the data to comprehensively assess LLM performance. To reduce human efforts in some redundant tasks, we proposed an automatic process to generate our dataset.
@@
    Term Extraction and Definition Generation: We extracted key terms from the questions and answers across all 57 subjects using the assistance of gpt-3.5-turbo. The process involved careful few-shot prompting, and we separately extracted the contexts of questions or answers alone to ensure the model focused on extracting terms rather than solving the questions. We also retrieved appropriate definitions within the specific subject for each extracted term. For terms where the automated process provided irrelevant or inaccurate definitions, we manually reviewed and corrected these entries.
@@
The MMLU-SR dataset was created using these meticulous steps. We formed both development and test sets, with the development set used for few-shot learning and the test set reserved for evaluation. This structured approach ensured that the dataset effectively tested the reasoning abilities of LLMs, differentiating between simple pattern recognition and genuine understanding. To effectively demonstrate how our MMLU-SR dataset can challenge more sophisticated models, we use Fig. 2, Fig. 3, and Fig. 4 that feature responses from ChatGPT-4. For comparative insights, example responses of ChatGPT-3.5 are available from Table 8, Table 9, and Table 10 in Appendix B.
Figure 2: Example ChatGPT-4 output of MMLU-SR “Question Only”.
Figure 3: Example ChatGPT-4 output of MMLU-SR “Answer Only”
@@
Fig. 2 illustrates an example where the dummy word ‘Hat’ is used to replace term ‘enzyme assay’ and is defined as “the name given to any laboratory technique that measures enzyme activity within a sample.” In this case, ChatGPT-4 successfully recognized ‘Hat’ and understood that the critical aspect was whether the diseases listed in the answer choices could be identified through enzyme activity. The model systematically eliminated each answer option, except for the last one, as it was the only choice that could be measured through enzyme activity tests. This setup effectively tests the model’s ability to comprehend and reason based on the provided definitions, rather than relying on pre-trained knowledge of the term ‘enzyme assay’.
@@
Our “Answer Only” dataset modifies the answer choices in most text from the original MMLU dataset similarly to the “Question Only” dataset and keeps the questions unchanged. We ensure that at least one answer choice is modified among all four choices. In specific cases, such as “moral scenarios” where the answers are simple TRUE or FALSE choices, we make substitutions like “Suppose ‘Jack’ means ‘True’.” and “Suppose ‘Luck’ means ‘False’.” Despite these modifications, Table 3 still shows the accuracy drops among all models compared to the original “moral scenarios” sheet from MMLU. This underscores the challenge posed to models in adapting to these symbolic substitutions. Fig. 3 demonstrates that ChatGPT-4 was able to recognize the replaced terms in answer choices A, B, and C, identifying ‘Bard’ as ‘Cystic Fibrosis’, ‘Jump’ as ‘Fragile X Syndrome’, and ‘Zelly’ as ‘Albinism’. The model incorrectly identified the term ‘Dummy’ as ‘Huntington’s Disease’, while the correct term is ‘Tay-Sachs Disease’. Both disorders are indeed genetic, but they are distinct in their genetic causes and manifestations. It appears that ChatGPT-4, focusing on the broader category of ‘genetic disorder’ from the provided definition, inadvertently linked the description to the wrong disease. Such misidentification led the model to persist in incorrectly affirming that choice A (‘Bard’ as ‘Cystic Fibrosis’) was the correct answer (it is not).
@@
Figure 4: Example ChatGPT-4 output of MMLU-SR “Question and Answer”.
Our “Question and Answer” dataset integrates elements from both the “Question Only” and “Answer Only” datasets, replacing fundamental terms in both the questions and answer choices with dummy words followed by their definitions. As illustrated in Fig. 4, ChatGPT-4 successfully interpreted the original terms for each replaced term in answer choices A through C.
However, similar to the results seen in Fig. 3, the model incorrectly recognized the term in the last answer choice D (‘Dummy’ for Huntington’s Disease), leading to an incorrect answer. This outcome contrasts with Fig. 2, where ChatGPT-4 correctly answered the questions when only the questions were modified. This illustrates that as complexity in context increases, with terms being replaced in both questions and answers, the model struggles to accurately identify the correct original term, consequently leading to an incorrect answer choice.

### v2

# MMLU-SR: A Benchmark for Stress-Testing Reasoning Capability of Large Language Models

@@ Abstract
MMLU-SR: A Benchmark for Stress-Testing
Reasoning Capability of Large Language Models
Wentian Wangthanks: Visiting student at Rutgers ML Lab. USC
Sarthak Jain Rutgers                        Paul Kantor Rutgers & UW-Madison                        Jacob Feldman Rutgers                        Lazaros Gallos Rutgers                        Hao Wang Rutgers
@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive quantitative performance on a wide range of benchmarks, natural language processing Zellers et al. (2019); Wang et al. (2019), general knowledge question-answeringHendrycks et al. (2021); Clark et al. (2018), and coding Chen et al. (2021); others (2021). Additionally, by integrating with some advanced prompting techniques, such as Chain-of-Thought (CoT) Wei et al. (2023) and its variants Yao et al. (2023); Trivedi et al. (2023); Zhang et al. (2023), LLMs seem to exhibit a certain level of reasoning abilities including mathematics Zhang et al. (2024) and even causal inference/discovery Vashishtha et al. (2023); Wang et al. (2020); Mao et al. (2022); Gupta et al. (2021). However, some studies Oren et al. (2023) have raised concerns about data leakage (i.e., training models on the test sets), potentially rendering these results unreliable. These seemingly contradictory findings prompt the question of whether LLMs are genuinely performing reasoning tasks or merely predicting the next token. If LLMs are truly capable of reasoning, they should remain unaffected by the replacement of key symbols within the test set.
A hallmark of human intelligence is the ability to handle abstract concepts and to associate them with arbitrary terms Penn et al. (2008). With a few exceptions such as onomatopoeia, the connection between particular words and particular meanings is arbitrary, and identical concepts are invoked by different words in different human languages (e.g. dog vs chien). Similarly, human reasoners are capable of analogizing structural relationships from one domain to another, meaning that conceptual equivalence can be retained even when details change Gentner and Medina (1998). It follows that true human-like comprehension should be unimpaired when terms are substituted for synonymous terms, as long as the substitution is comprehensibly defined.
We wondered whether LLM peformance reflects true human-like comprehension in this sense, or whether it relies heavily on the specific terms used on training corpora. To assess this, we propose MMLU-SR, a new benchmark dataset that uses symbol replacement to remove some important terms from the questions and answers as shown in Figure 1. Instead of relying on memorized terms, this approach tests whether LLMs can reason using the definitions and concepts of those terms, ensuring a more robust evaluation of their understanding.
Our evaluations on GPT-3.5/4, Gemini, and Llama3 families showed significantly lower performance on MMLU-SR compared to the original MMLU, demonstrating the effectiveness of our approach in preventing models from exploiting memorized data. MMLU-SR thus provides a more challenging and revealing test of LLMs’ true reasoning abilities and understanding.
@@
2 Related Works
MMLU Variants Benchmarks. MMLU Variants such as CMMLU Li et al. (2024) and TMMLU+ Tam et al. (2024) are adaptations of the MMLU benchmark for non-English languages; they translate the original MMLU questions and answers into other languages, providing a way to evaluate language models’ performance in non-English contexts. These benchmarks are crucial for assessing the generalizability and robustness of models across different languages and cultural settings. They preserve the original structure and intent of MMLU while enabling a broader examination of multilingual capabilities.
Reasoning Capabilities Benchmarks. Several advanced reasoning benchmarks have been developed to evaluate the reasoning capabilities of language models. AGIEval Zhong et al. (2023) includes standardized tests and civil service exams to assess reasoning and problem-solving skills in academic and professional scenarios. BoolQ Clark et al. (2019) comprises over 15,000 real yes/no questions paired with Wikipedia passages to test the ability of models to infer answers from contextual information. GSM8K Cobbe et al. (2021) features 8.5K grade-school math problems requiring multistep operations, targeting the evaluation of basic to intermediate mathematical problem-solving abilities. DROP Dua et al. (2019), an adversarially-created reading comprehension benchmark, challenges models to navigate references and perform discrete operations such as addition and sorting, thus evaluating their capacity to understand complex texts and execute logical reasoning tasks. Beyond purely language-based evaluation, on the multimodal front, MMNeedle Wang et al. (2024) introduced one of the first multimodal benchmarks to evaluate long-context multimodal reasoning capabilities of multimodal LLMs.
@@ 3 MMLU-SR Dataset
Figure 2: Example ChatGPT-4 output of MMLU-SR ‘Question Only”.
@@ 3.1 Dataset Construction
We have developed the MMLU-SR benchmark to rigorously evaluate the reasoning and understanding capabilities of LLMs. Inspired by ObjectNet Borji (2020), our benchmark contains three subsets: “Question Only”, “Answer Only”, and “Question and Answer”, each offering a unique perspective on the data to comprehensively assess LLM performance. To reduce human efforts in some redundant tasks, we proposed an automatic process to generate our dataset.
@@
    Term Extraction and Definition Generation: We extracted key terms from the questions and answers across all 57 subjects using the assistance of gpt-3.5-turbo. The process involved careful few-shot prompting, and we separately extracted the contexts of questions or answers alone to ensure the model focused on extracting terms rather than solving the questions. We also retrieved appropriate definitions within the specific subject for each extracted term. For terms where the automated process provided irrelevant or inaccurate definitions, we manually reviewed and corrected these entries (see Appendix F for details on the extent of manual modifications).
@@
The MMLU-SR dataset was created using these meticulous steps. We formed both development and test sets, with the development set used for few-shot learning and the test set reserved for evaluation. This structured approach ensured that the dataset effectively tested the reasoning abilities of LLMs, differentiating between simple pattern recognition and genuine understanding. To effectively demonstrate how our MMLU-SR dataset can challenge more sophisticated models, we use Figure 2, Figure 3, and Figure 4 that feature responses from ChatGPT-4. For comparative insights, example responses of ChatGPT-3.5 are available from Table 8, Table 9, and Table 10 in Appendix B.
@@
Figure 3: Example ChatGPT-4 output of MMLU-SR “Answer Only”.
Figure 2 illustrates an example where the dummy word ‘Hat’ is used to replace term ‘enzyme assay’ and is defined as “the name given to any laboratory technique that measures enzyme activity within a sample.” In this case, ChatGPT-4 successfully recognized ‘Hat’ and understood that the critical aspect was whether the diseases listed in the answer choices could be identified through enzyme activity. The model systematically eliminated each answer option, except for the last one, as it was the only choice that could be measured through enzyme activity tests. This setup effectively tests the model’s ability to comprehend and reason based on the provided definitions, rather than relying on pre-trained knowledge of the term ‘enzyme assay’.
Table 1: Performance of gpt-4o-mini,gpt-4o, gemini-1.5-pro, and llama3-70b.
Dataset Humanities Social Sciences STEM Other Average
GPT-4o-mini
MMLU (5-shot) 0.793 0.858 0.689 0.782 0.771
Question Only (5-shot) 0.744 0.792 0.621 0.724 0.710
Answer Only (5-shot) 0.659 0.738 0.602 0.651 0.655
Question and Answer (5-shot) 0.588 0.666 0.531 0.585 0.585
GPT-4o
MMLU (5-shot) 0.880 0.906 0.771 0.854 0.845
Question Only (5-shot) 0.838 0.856 0.702 0.811 0.792
Answer Only (5-shot) 0.764 0.824 0.705 0.760 0.757
Question and Answer (5-shot) 0.708 0.754 0.635 0.712 0.695
Gemini-1.5-pro
MMLU (5-shot) 0.849 0.881 0.802 0.815 0.832
Question Only (5-shot) 0.795 0.836 0.700 0.754 0.764
Answer Only (5-shot) 0.741 0.816 0.747 0.739 0.758
Question and Answer (5-shot) 0.690 0.752 0.670 0.681 0.694
Llama3-70B
MMLU (5-shot) 0.681 0.868 0.697 0.814 0.765
Question Only (5-shot) 0.635 0.812 0.631 0.770 0.712
Answer Only (5-shot) 0.539 0.683 0.565 0.622 0.602
Question and Answer (5-shot) 0.523 0.653 0.536 0.591 0.576
Dataset Humanities Social Sciences STEM Other Average
GPT-4o-mini
Question Only (5-shot) 6.18% 7.69% 9.87% 7.42% 7.91%
Answer Only (5-shot) 16.90% 13.99% 12.63% 16.75% 15.05%
Question and Answer (5-shot) 25.85% 22.38% 22.93% 25.19% 24.12%
GPT-4o
Question Only (5-shot) 4.77% 5.52% 8.95% 5.03% 6.27%
Answer Only (5-shot) 13.18% 9.05% 8.56% 11.01% 10.41%
Question and Answer (5-shot) 19.55% 16.78% 17.64% 16.63% 17.75%
Gemini-1.5-pro
Question Only (5-shot) 6.36% 5.11% 12.72% 7.48% 8.17%
Answer Only (5-shot) 12.72% 7.38% 6.86% 9.33% 8.89%
Question and Answer (5-shot) 18.73% 14.64% 16.46% 16.44% 16.59%
Llama3-70B
Question Only (5-shot) 6.75% 6.45% 9.47% 5.41% 6.93%
Answer Only (5-shot) 20.85% 21.31% 18.94% 23.59% 21.31%
Question and Answer (5-shot) 23.20% 24.77% 23.10% 27.40% 24.71%
@@
Figure 4: Example ChatGPT-4 output of MMLU-SR “Question and Answer”.
Our “Answer Only” dataset modifies the answer choices in most text from the original MMLU dataset similarly to the “Question Only” dataset and keeps the questions unchanged. We ensure that at least one answer choice is modified among all four choices. In specific cases, such as “moral scenarios” where the answers are simple TRUE or FALSE choices, we make substitutions like “Suppose ‘Jack’ means ‘True’.” and “Suppose ‘Luck’ means ‘False’.” Despite these modifications, Table 3 still shows the accuracy drops among all models compared to the original “moral scenarios” sheet from MMLU. This underscores the challenge posed to models in adapting to these symbolic substitutions. Figure 3 demonstrates that ChatGPT-4 was able to recognize the replaced terms in answer choices A, B, and C, identifying ‘Bard’ as ‘Cystic Fibrosis’, ‘Jump’ as ‘Fragile X Syndrome’, and ‘Zelly’ as ‘Albinism’. The model incorrectly identified the term ‘Dummy’ as ‘Huntington’s Disease’, while the correct term is ‘Tay-Sachs Disease’. Both disorders are indeed genetic, but they are distinct in their genetic causes and manifestations. It appears that ChatGPT-4, focusing on the broader category of ‘genetic disorder’ from the provided definition, inadvertently linked the description to the wrong disease. Such misidentification led the model to persist in incorrectly affirming that choice A (‘Bard’ as ‘Cystic Fibrosis’) was the correct answer (it is not).
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Formal Logic 0.730 0.698 0.532 0.603 0.500 0.484 0.643 0.579 0.516 0.556 0.500 0.460 Logical Fallacies 0.902 0.902 0.853 0.883 0.834 0.810 0.853 0.847 0.663 0.834 0.841 0.564 Moral Disputes 0.882 0.832 0.847 0.832 0.806 0.769 0.777 0.830 0.630 0.711 0.749 0.653 Moral Scenarios 0.813 0.760 0.318 0.830 0.774 0.289 0.143 0.199 0.318 0.177 0.167 0.253 Philosophy 0.891 0.865 0.865 0.778 0.724 0.772 0.698 0.756 0.598 0.582 0.611 0.582 World Religions 0.901 0.895 0.906 0.895 0.836 0.895 0.842 0.813 0.696 0.825 0.772 0.684 High School European History 0.903 0.885 0.848 0.885 0.855 0.830 0.897 0.849 0.721 0.861 0.818 0.739 High School Us History 0.946 0.922 0.946 0.917 0.902 0.887 0.897 0.863 0.799 0.863 0.819 0.799 High School World History 0.937 0.920 0.945 0.924 0.920 0.916 0.907 0.865 0.806 0.882 0.827 0.840 Prehistory 0.948 0.901 0.910 0.904 0.836 0.793 0.843 0.803 0.670 0.790 0.769 0.670 International Law 0.942 0.926 0.868 0.901 0.860 0.868 0.934 0.843 0.769 0.835 0.802 0.760 Jurisprudence 0.898 0.861 0.852 0.852 0.861 0.806 0.861 0.806 0.602 0.722 0.750 0.556 Professional Law 0.749 0.666 0.616 0.683 0.627 0.583 0.641 0.585 0.461 0.563 0.544 0.461
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Econometrics 0.711 0.702 0.693 0.588 0.579 0.570 0.640 0.614 0.561 0.535 0.535 0.421 High School Macroeconomics 0.921 0.880 0.821 0.849 0.785 0.779 0.813 0.785 0.628 0.721 0.715 0.572 High School Microeconomics 0.971 0.929 0.870 0.903 0.870 0.773 0.857 0.815 0.664 0.769 0.744 0.571 High School Government And Politics 0.984 0.974 0.969 0.979 0.943 0.938 0.943 0.922 0.798 0.922 0.845 0.782 Public Relations 0.836 0.746 0.755 0.755 0.755 0.736 0.664 0.682 0.600 0.627 0.646 0.555 Security Studies 0.824 0.841 0.824 0.788 0.792 0.767 0.731 0.796 0.673 0.633 0.714 0.624 Us Foreign Policy 0.930 0.940 0.930 0.920 0.930 0.890 0.870 0.880 0.740 0.810 0.810 0.780 Human Sexuality 0.931 0.893 0.855 0.924 0.855 0.840 0.863 0.847 0.710 0.802 0.756 0.756 Sociology 0.935 0.891 0.920 0.900 0.896 0.841 0.881 0.881 0.806 0.831 0.851 0.786 High School Geography 0.955 0.939 0.924 0.894 0.909 0.833 0.884 0.864 0.737 0.813 0.813 0.662 High School Psychology 0.965 0.938 0.921 0.923 0.917 0.884 0.927 0.912 0.719 0.872 0.859 0.739 Professional Psychology 0.908 0.895 0.845 0.845 0.801 0.788 0.817 0.791 0.627 0.719 0.737 0.601
@@
Our “Question and Answer” dataset integrates elements from both the “Question Only” and “Answer Only” datasets, replacing fundamental terms in both the questions and answer choices with dummy words followed by their definitions. As illustrated in Figure 4, ChatGPT-4 successfully interpreted the original terms for each replaced term in answer choices A through C. However, similar to the results seen in Figure 3, the model incorrectly recognized the term in the last answer choice D (‘Dummy’ for Huntington’s Disease), leading to an incorrect answer. This outcome contrasts with Figure 2, where ChatGPT-4 correctly answered the questions when only the questions were modified. This illustrates that as complexity in context increases, with terms being replaced in both questions and answers, the model struggles to accurately identify the correct original term, consequently leading to an incorrect answer choice.
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Abstract Algebra 0.660 0.690 0.380 0.470 0.550 0.370 0.640 0.730 0.400 0.460 0.520 0.400 College Mathematics 0.490 0.680 0.510 0.420 0.630 0.490 0.440 0.650 0.460 0.410 0.610 0.480 High School Statistics 0.769 0.866 0.699 0.708 0.708 0.657 0.750 0.829 0.620 0.644 0.662 0.597 Elementary Mathematics 0.735 0.921 0.606 0.675 0.786 0.521 0.706 0.900 0.561 0.661 0.825 0.497 High School Mathematics 0.541 0.700 0.422 0.537 0.504 0.356 0.541 0.615 0.426 0.511 0.526 0.367 Astronomy 0.947 0.901 0.921 0.908 0.829 0.849 0.888 0.849 0.697 0.855 0.796 0.684 College Physics 0.686 0.716 0.559 0.559 0.647 0.451 0.618 0.745 0.431 0.480 0.608 0.422 Conceptual Physics 0.911 0.932 0.783 0.804 0.757 0.677 0.791 0.843 0.494 0.685 0.698 0.447 High School Physics 0.748 0.782 0.563 0.649 0.556 0.530 0.589 0.616 0.477 0.543 0.596 0.450 College Chemistry 0.570 0.610 0.580 0.540 0.550 0.570 0.550 0.530 0.480 0.480 0.560 0.470 High School Chemistry 0.759 0.788 0.734 0.709 0.685 0.631 0.670 0.680 0.537 0.586 0.626 0.468 College Biology 0.951 0.868 0.931 0.938 0.882 0.854 0.924 0.861 0.708 0.833 0.826 0.625 High School Biology 0.958 0.929 0.903 0.932 0.893 0.858 0.884 0.858 0.713 0.858 0.829 0.729 College Computer Science 0.790 0.790 0.670 0.690 0.610 0.650 0.760 0.730 0.610 0.670 0.660 0.570 Computer Security 0.840 0.820 0.830 0.830 0.770 0.750 0.760 0.730 0.660 0.760 0.610 0.720 High School Computer Science 0.910 0.920 0.870 0.860 0.880 0.790 0.880 0.910 0.820 0.850 0.870 0.740 Machine Learning 0.777 0.714 0.652 0.661 0.643 0.589 0.643 0.661 0.527 0.580 0.580 0.509 Electrical Engineering 0.841 0.807 0.745 0.752 0.724 0.655 0.655 0.710 0.510 0.566 0.655 0.490
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Anatomy 0.911 0.793 0.807 0.874 0.733 0.726 0.815 0.667 0.563 0.726 0.659 0.578 Clinical Knowledge 0.898 0.838 0.849 0.811 0.785 0.740 0.796 0.755 0.638 0.713 0.709 0.608 College Medicine 0.832 0.844 0.757 0.780 0.786 0.740 0.798 0.763 0.647 0.717 0.740 0.659 Human Aging 0.830 0.807 0.807 0.794 0.744 0.758 0.704 0.740 0.457 0.632 0.691 0.471 Medical Genetics 0.960 0.910 0.830 0.900 0.850 0.820 0.840 0.780 0.570 0.830 0.740 0.550 Nutrition 0.899 0.876 0.853 0.863 0.758 0.804 0.798 0.784 0.663 0.699 0.703 0.647 Professional Medicine 0.956 0.864 0.868 0.919 0.776 0.868 0.901 0.783 0.754 0.842 0.735 0.754 Virology 0.578 0.578 0.536 0.548 0.506 0.488 0.524 0.542 0.452 0.524 0.494 0.404 Business Ethics 0.860 0.850 0.750 0.890 0.780 0.720 0.750 0.670 0.500 0.710 0.640 0.480 Management 0.913 0.893 0.913 0.883 0.816 0.903 0.757 0.835 0.728 0.767 0.767 0.650 Marketing 0.949 0.940 0.923 0.906 0.927 0.880 0.838 0.846 0.615 0.808 0.803 0.662 Global Facts 0.650 0.600 0.530 0.540 0.540 0.430 0.580 0.690 0.540 0.520 0.470 0.410 Miscellaneous 0.955 0.955 0.903 0.932 0.877 0.860 0.861 0.847 0.692 0.840 0.791 0.616 Professional Accounting 0.766 0.663 0.638 0.716 0.674 0.596 0.681 0.638 0.514 0.631 0.596 0.489